Need help:
One card is selected from a deck of cards. Find the probability of selecting a black card or a queen.
A) 1/26
B) 15/26
C) 7/13
D) 27/52

Answer by Alan3354(69443)

One card is selected from a deck of cards. Find the probability of selecting a black card or a queen.
-------------------
Black card = 1/2
Queen = 4/52 = 1/13
It's "or", so you add them.
1/2 + 1/13 = 15/26